# Correctness Assessment of Code Generated by Large Language Models Using Internal Representations

### v1

@@ 4.3 Experimental Procedure
Baselines: We evaluate the performance of Openia by comparing it with two baselines:
@@
  • Classification-based [37, 38, 39]: These approaches focus on capturing the semantics of code to predict its correctness. Specifically, pre-trained models such as CodeBERT [40] and CodeT5 [41]/CodeT5+ [42] are employed to encode the semantics of the code, which is generated by the LLM. The resulting embeddings are then used to train a classifier which determines whether the given code is correct.
  • Prompting-based [43]: These approaches leverage the reasoning and generalization capabilities of LLMs to access the code correctness without requiring additional fine-tuning or training. Specifically, prompting techniques such as zero-shot and few-shot learning are applied to directly query the LLM for predictions on the correctness of the given code.
@@
For a fair comparison, in all classification-based approaches and Openia, we apply the same neural architecture for the classifier containing the input, output layers, and two hidden layers with 128 and 64 neurons correspondingly. And the training process runs for 50 epochs with a batch size of 32, using a learning rate of 0.001.
@@
Procedure: For comparison with the baselines, the internal state of the last generated token in the last layer of the studied LLM is extracted for each query as a representation to assess the correctness of the corresponding generated output. This choice is motivated by the fact that the internal state from the last layer is closest to the model’s final decision, making it a logical and effective representation for evaluating the correctness of the generated code. Moreover, the last token is generated based on both the prompt input and the entire generated output, allowing its internal state to encapsulate the most comprehensive information processed by the model. Therefore, this state is selected for conducting experiments to answer this research question. While this experiment focuses on leveraging the last token’s internal state for comparisons with baselines, the impacts of other tokens and layers are investigated separately in RQ2.
@@
In this experiment, we compare the performance of Openia with the baselines across multiple settings.
@@
    • Cross-task: The tasks within a benchmark are split in an 9:1 ratio, where 90% of the tasks are used to extract internal representations for training the probing classifier, and 10% of the remaining tasks are used for testing.
  • Repo(sitory)-level code generation: This evaluates the correctness of code generated for existing projects. Unlike independent-unit code generation, this setting involves understanding the broader context of the codebase [44, 45]. This setting is particularly challenging, as it requires LLMs to maintain contextual consistency and correctness across multiple units of code within a project.
RQ2. Layer and token selection analysis: This experiment investigates the impact of internal states from different layers and token positions on Openia’s performance in assessing the correctness of generated code. To answer this research question, we systematically evaluate the performance of Openia when it is trained and tested using internal states extracted from various layers and token positions across multiple settings.
RQ3: Sensitivity analysis: This experiment examines Openia’s ability to generalize across programming languages. Specifically, we evaluate whether Openia, trained on the internal states from the generation of code in several programming languages, can effectively predict the correctness of code in a completely different language. In addition, we also investigate the performance of Openia with code generated for tasks of different difficulty levels.
@@ 4.4 Metrics
The task of accessing the correctness of code generated by LLM can be formulated as a binary classification task. Specifically, given a generated code snippet (along with related information, such as its internal representations), Openia and the baseline approaches aim to predict whether the code is correct or not. To evaluate the performance of the approaches, we adapt widely used classification metrics: Accuracy, Precision, Recall, and F1-Score.
Given the potential class imbalance in the dataset, as highlighted in Table I, where the ratio of correct to incorrect code snippets may vary significantly, it is critical to ensure a fair evaluation across both classes. To address this, we employ weighted metrics, including Weighted Accuracy, Weighted Prediction, Weighted Recall, and Weighted F1-score. These metrics compare the corresponding result for each label and return the average considering the proportion for each label in the test set.
@@ 5.1.1 Performance Comparison
TABLE II: Correctness assessment performance for independent-unit code generation
@@
DeepSeek-1.3B Cross-benchmark Zeroshot 0.55 0.62 0.55 0.50
@@
Cross-task Zeroshot 0.42 0.42 0.42 0.42
@@
DeepSeek-6.7B Cross-benchmark Zeroshot 0.32 0.69 0.32 0.26
@@
Cross-task Zeroshot 0.68 0.78 0.68 0.57
@@
CodeLlama-7B Cross-benchmark Zeroshot 0.61 0.37 0.61 0.46
@@
Cross-task Zeroshot 0.50 0.25 0.50 0.33
@@
Openia 0.69 0.76 0.75 0.75
MagicCoder-7B Cross-benchmark Zeroshot 0.27 0.07 0.27 0.11
@@
Cross-task Zeroshot 0.48 0.73 0.48 0.47
@@
Table II and Table III show the performance of correctness assessment for independent-unit code generation and repo-level code generation for the studied code LLMs, DeepSeekCoder 1.3B and 6.7B, CodeLlama-7B, and MagicCoder-7B. Overall, Openia consistently outperforms all baselines across all experimental settings in all evaluation metrics and for all studied code LLMs, demonstrating its robustness in diverse scenarios.
For independent-unit code generation shown in Table II, in Cross-benchmark, Openia achieves an accuracy of up to 0.69 (for CodeLlama-7B), surpassing the baseline approaches by from 9% to 118% relatively. In Cross-task, Openia obtains accuracies of 0.69 (for CodeLlama-7B) to 0.79 (for DeepSeek-6.7B), exceeding the baseline performance by up to 2 times. This superior performance highlights the effectiveness of Openia by leveraging the internal representations of the LLM in assessing the correctness of generated code.
For repo-level code generation shown in Table III, Openia also consistently surpasses the baseline approaches. For example, Openia obtains the F1-Scores from 0.73 to 0.79, while the corresponding figures of Zeroshot, CodeBERT, and CodeT5+ range from 0.50 to 0.71. Especially for CodeLlama-7B, Openia’s accuracy is 0.81, and the improvements by Openia in F1-Score are from 11% to 46%. These results highlight the effectiveness of leveraging internal representations of the LLM, making Openia more reliable in assessing the correctness of the real-world repo-level code.
In our experiments, Zeroshot, which prompts the LLM to determine whether a given code snippet is correct, exhibits an imbalance in handling positive and negative cases. Specifically, while this approach can achieve competitive Precision, it struggles with Recall, especially for DeepSeek Coder variants. For DeepSeek-6.7B, Zeroshot obtains a Precision of 0.69 and 0.70 in independent-unit and repo-level code generation settings, respectively, which is comparable to the other approaches, yet its Recall is up to about two times lower than the others. The reason is that the LLM tends to be over-confident in the correctness of its generated code, frequently labeling its generated code as “correct” even when it is actually incorrect.
Traditional classification-based approaches, which predict code correctness using the code semantics encoded by pre-trained models such as CodeBERT and CodeT5/CodeT5+, demonstrate a significantly stronger performance compared to Zeroshot. Despite their strengths, however, these baselines are outperformed by Openia, as evidenced by consistently higher accuracy, precision, recall, and F1-Scores across all models in Table II and Table III. For repo-level code generation, Openia achieves an F1-Score of 0.73 with DeepSeek-6.7B, compared to 0.69 for CodeT5+. Especially for CodeLlama-7B, the margin of improvement is even more pronounced, with Openia attaining F1-Scores of 0.79, respectively, compared to 0.71 and 0.69 for the classifier based on CodeT5+ and CodeBERT. These results suggest that relying on pre-trained models to encode generated code, while effective to an extent, is insufficient for capturing the nuanced, dynamic, and process-oriented aspects of code correctness.
Unlike these black-box (closed-box) approaches that rely solely on final outputs, Openia is a white-box (open-box) approach that examines the internal workings of the LLM to make predictions. By analyzing the intermediate states of the LLM during code generation, Openia assesses code correctness based on a more comprehensive set of signals derived from these states . This approach uncovers hidden patterns associated with the generated code, mitigating the overconfidence often seen in the Zeroshot approach and providing a more nuanced and reliable evaluation. Moreover, the internal representations utilized by Openia not only encode the code features as CodeBERT and CodeT5+ but also capture the model’s reasoning process. This reasoning process information reflects not just dataset-specific patterns but also fundamental programming underlying programming principles. Openia can dynamically adapt to varying requirements, programming constructs, and output expectations across different tasks and benchmarks, enhancing its generality and robustness. This adaptability enhances its generality and robustness, enabling it to consistently outperform baseline methods across all experimental settings. Additionally, by leveraging these in-process signals, Openia provides a unique advantage: it can detect potential issues and assess correctness as the code is being generated, enabling real-time quality assurance.
TABLE III: Correctness assessment performance for repo-level code generation
@@
DeepSeek-1.3B ZeroShot 0.24 0.76 0.24 0.27
@@
DeepSeek-6.7B ZeroShot 0.48 0.70 0.48 0.50
@@
MagicCoder-7B ZeroShot 0.74 0.64 0.74 0.65
@@
CodeLlama-7B ZeroShot 0.51 0.65 0.51 0.54
CodeBERT 0.74 0.68 0.74 0.69
CodeT5+ 0.76 0.71 0.76 0.71
Openia 0.81 0.79 0.81 0.79
@@ 5.1.2 Code length analysis
Fig. 2 illustrates that Openia achieves its best performance with short generated code. Specifically, Openia accurately predicts the correctness of the code snippets with lengths ranging from 1 to 30 tokens, achieving an F1-Score of 0.81. This is because short code snippets are often simple and contain few dependencies. Thus, the internal states of the LLM the semantic and syntactic features of the code can be fully encoded in the internal states of the LLMs, thereby facilitating accurate predictions.
However, the F1-score of Openia decreases as the code length increases, reaching its lowest point (0.66) for medium-length code containing 51-70 tokens. Longer code tends to be more complex, making it challenging for the internal states to fully capture all critical information. This limitation results in a performance drop for Openia.
Interestingly, Openia exhibits performance recovery for longer code, accurately predicting code correctness of exceeding 70 tokens. The reason is that the richer context and redundancy in longer code allow the LLM to better resolve dependencies, reinforce patterns, and effectively focus on relevant sections, leading to improved prediction accuracy. These results highlight the adaptability of Openia in leveraging the LLM’s representations across varying lengths.

### v2

@@ 4.3 Experimental Procedure
Baselines: We evaluated the performance of Openia by comparing it with three baseline approaches:
  • Inference-time representation-based methods [37, 38, 39, 40]: These approaches leverage the internal representations of LLMs during inference to access the correctness of generated outputs. To the best of our knowledge, no prior work has specifically explored internal representations for evaluating the correctness of code generated by LLMs. For comparative evaluation, we adopt LLM-Check [40], a method originally proposed for assessing the quality of LLM-generated outputs in natural language question answering tasks. Specifically, LLM-Check utilizes the diversified scoring methods, including internal attention kernel maps, hidden activations and output prediction probabilities, from different model components to maximize the capture of hallucinations.
@@
  • Post-hoc classification-based methods [41, 42, 43]: These approaches focus on capturing the semantics of generated code to predict its correctness. Specifically, pre-trained models such as CodeBERT [44] and CodeT5 [45] or CodeT5+ [46] are employed to encode the semantics of the LLM-generated code. A classifier is then trained on these embeddings to determine whether the given code is correct or not.
  • LLM-as-A-Judge (LLM-AJ) [37]: These approaches leverage the inherent reasoning and generalization capabilities of LLMs to access code correctness without requiring additional fine-tuning or training. Following the setup of prior work [37], we adopt a zero-shot prompting strategy using the same prompt design, where the model is directly queried regarding the correctness of a given code snippet. The full prompt details are available on our website [36]. In our experiments, we employ zero-shot inference with both the model under study (In-house LLM-AJ) (e.g., DeepSeek Coder, Code Llama, or Magicoder) and an external reasoning model o4-mini (External LLM-AJ), which serves as an independent verifier to evaluate code correctness.
@@
For a fair comparison, we applied the same neural architecture for the classification models in both post-hoc methods and Openia. In each approach, the classifier contains the input, output layers, and two hidden layers with 128 and 64 neurons correspondingly. The training is conducted for 50 epochs with a batch size of 32 and a learning rate of $10^{-3}$.
@@
Procedure: To compare Openia’s performance with the baselines, we extracted the internal state of the last generated token from the last layer of the studied LLM for each query as a representation to assess the correctness of the corresponding generated output. This choice is motivated by the fact that the internal state from the last layer is closest to the model’s final decision, making it a logical and effective representation for evaluating the correctness of the generated code. Moreover, the last token is generated based on both the prompt input and the entire generated output, allowing its internal state to encapsulate the most comprehensive information processed by the model. Therefore, this state is selected for conducting experiments to answer this research question. Note that while this experiment focuses on leveraging the last token’s internal state for comparison with baselines, the impacts of other tokens and layers are investigated separately in RQ2.
Note that in our comparative study, while both Openia and the post-hoc classification-based methods provide their code quality assessment after the code generation is complete, they are fundamentally different in terms of the input. The post-hoc classification-based methods rely exclusively on the final output code produced after the generation process. In contrast, Openia gives its assessment based on the code LLMs’ internal states which are the intermediate artifacts produced during the generation process. This design also enables Openia to support early code correctness assessment. For example, by analyzing the internal representations of early generated tokens/layers, correctness assessments could be made even before code generation finishes, allowing proactive interventions.
@@
In this experiment, we compared the performance of Openia with the baselines across multiple settings.
@@
    • Cross-task: The tasks within a benchmark are split in a 9:1 ratio, where 90% of the tasks were used to extract internal representations for training and validating the probing classifier, and 10% of the remaining tasks are used for testing.
  • Repo(sitory)-level code generation: This evaluates the correctness of code generated for existing projects. Unlike independent-unit code generation, repo-level tasks require LLMs to maintain contextual consistency across multiple code units, making them significantly more challenging [47, 48]. This setting measures how well Openia can handle larger, context-dependent codebases.
RQ2. Layer and token selection analysis: This experiment investigates the impact of internal representations from different layers and token positions on Openia’s performance in assessing code correctness. To answer this research question, we systematically evaluated the performance of Openia when it is trained and tested on internal states extracted from:
  • Different layers within the code LLM, ranging from shallow to deep layers, to determine how abstraction levels influence correctness prediction.
  • Various token positions, including first and last generated token, first and last generated code token, to understand their respective contribution to accuracy.
RQ3: Sensitivity analysis: This experiment examines Openia’s ability to generalize across programming languages. Specifically, we evaluated whether Openia, trained on the internal states extracted from the generation of code in several programming languages, can effectively predict the correctness of code in a completely different language. In addition, we also investigated the performance of Openia with code generated for tasks of different difficulty levels.
@@ 4.4 Metrics
The task of accessing the correctness of code generated by LLM can be formulated as a binary classification task. Specifically, given a generated code snippet (along with related information, such as its internal representations), Openia and the baseline approaches aim to predict whether the code is correct or not. To evaluate the performance of the approaches, we adapted widely used classification metrics: Accuracy, Precision, Recall, and F1-Score.
Given the potential class imbalance in the dataset, as highlighted in Table 1, where the ratio of correct to incorrect code snippets may vary significantly, it is critical to ensure a fair evaluation across both labels. To address this, we employed weighted metrics, including Weighted Accuracy, Weighted Precision, Weighted Recall, and Weighted F1-score. These metrics compare the corresponding result for each label and return the average considering the proportion for each label in the test set.
@@ 5.1.1 Performance Comparison
Table 2: Correctness assessment performance for Independent-unit Code Generation
@@
DeepSeek Coder-1.3B Cross-benchmark In-house LLM-AJ 0.55 0.62 0.55 0.50
External LLM-AJ 0.93 0.93 0.93 0.93
@@
Cross-task In-house LLM-AJ 0.42 0.42 0.42 0.42
External LLM-AJ 0.86 0.86 0.86 0.86
@@
DeepSeek Coder-6.7B Cross-benchmark In-house LLM-AJ 0.32 0.69 0.32 0.26
External LLM-AJ 0.94 0.94 0.94 0.94
@@
Cross-task In-house LLM-AJ 0.68 0.78 0.68 0.57
External LLM-AJ 0.75 0.77 0.75 0.75
@@
Code Llama-7B Cross-benchmark In-house LLM-AJ 0.61 0.37 0.61 0.46
External LLM-AJ 0.97 0.97 0.97 0.97
@@
Cross-task In-house LLM-AJ 0.50 0.25 0.50 0.33
External LLM-AJ 0.75 0.75 0.75 0.75
@@
LLM-Check 0.66 0.45 0.66 0.53
Openia 0.75 0.76 0.75 0.75
Code Llama-13B Cross-benchmark In-house LLM-AJ 0.41 0.16 0.41 0.23
External LLM-AJ 0.95 0.95 0.95 0.95
CodeBERT 0.61 0.61 0.61 0.61
CodeT5+ 0.62 0.62 0.62 0.62
LLM-Check 0.56 0.65 0.56 0.56
Openia 0.64 0.65 0.64 0.64
Cross-task In-house LLM-AJ 0.38 0.14 0.38 0.21
External LLM-AJ 0.75 0.75 0.75 0.75
CodeBERT 0.65 0.66 0.65 0.65
CodeT5+ 0.67 0.67 0.67 0.67
LLM-Check 0.64 0.68 0.64 0.57
Openia 0.78 0.79 0.78 0.78
Magicoder-7B Cross-benchmark In-house LLM-AJ 0.27 0.07 0.27 0.11
External LLM-AJ 0.94 0.95 0.94 0.94
@@
Cross-task In-house LLM-AJ 0.48 0.73 0.48 0.47
External LLM-AJ 0.80 0.81 0.80 0.80
@@
Table 3: Correctness assessment performance for Repo-level Code Generation
@@
DeepSeek Coder-1.3B In-house LLM-AJ 0.24 0.76 0.24 0.27
External LLM-AJ 0.87 0.87 0.87 0.87
@@
DeepSeek Coder-6.7B In-house LLM-AJ 0.48 0.70 0.48 0.50
External LLM-AJ 0.70 0.73 0.70 0.71
@@
Code Llama-7B In-house LLM-AJ 0.51 0.65 0.51 0.54
External LLM-AJ 0.73 0.76 0.73 0.74
CodeBERT 0.74 0.68 0.74 0.69
CodeT5+ 0.76 0.71 0.76 0.71
LLM-Check 0.77 0.59 0.77 0.67
Openia 0.81 0.79 0.81 0.79
Code Llama-13B In-house LLM-AJ 0.19 0.04 0.19 0.06
External LLM-AJ 0.77 0.79 0.77 0.78
CodeBERT 0.80 0.73 0.80 0.74
CodeT5+ 0.78 0.70 0.78 0.73
LLM-Check 0.80 0.66 0.80 0.73
Openia 0.81 0.77 0.81 0.78
Magicoder-7B In-house LLM-AJ 0.74 0.64 0.74 0.65
External LLM-AJ 0.70 0.74 0.70 0.72
@@
Table 2 and Table 3 show the performance of code correctness assessment for independent-unit code generation and repo-level code generation across the studied code LLMs, DeepSeek Coder-1.3B and 6.7B, Code Llama 7B and 13B, and Magicoder-7B. The evaluation includes LLM-as-a-judge (LLM-AJ) methods using the LLM under study (In-house LLM-AJ) and a reasoning LLM o4-mini (External LLM-AJ), post-hoc methods with CodeBERT and CodeT5+, and our approach Openia. Overall, Openia demonstrates robust and strong performance across all metrics and evaluation settings, maintaining its top results regardless of model size or task complexity.
In Cross-benchmark (Table 2), although Openia lags behind External LLM-AJ, it significantly outperforms In-house LLM-AJ and the other baseline methods. For example, Openia achieves an Accuracy of 0.69 (for Code Llama-7B), surpassing the the traditional post-hoc classification methods by a relative improvement of 10% to 18%. In Cross-task, Openia and External LLM-AJ achieve competitive performance. Openia obtains accuracies ranging from 0.73 (for DeepSeek Coder-1.3B) to 0.79 (for DeepSeek Coder-6.7B), significantly exceeding the performance of the post-hoc classification methods by up to 23%. This superior performance highlights the effectiveness of leveraging the internal representations of the LLM, demonstrating Openia’s capability to assess the correctness of generated code with high reliability.
For repo-level code generation (Table 3), Openia also demonstrate impressive performance, consistently surpasses the baseline approaches, including all the studied LLM-AJ methods, post-hoc methods, and LLM-Check. For example, Openia obtains F1-Scores ranging from 0.73 to 0.83, while the corresponding figures of In-house LLM-AJ and the post-hoc methods based on CodeBERT and CodeT5+ are between 0.27 and 0.83. Especially for DeepSeek Coder-1.3B, Openia achieves an Accuracy of 0.86 with an F1-Score improvement of up to three times over the baselines, while for Code Llama-7B, Openia’s Accuracy is 10% relatively better than that of External LLM-AJ. These findings underscore the advantage of leveraging internal representations, making Openia more reliable in assessing the correctness of the real-world and complex repo-level code generated by the LLMs.
In our experiments, for In-house LLM-AJ with the studied LLMs (i.e., DeepSeek Coder, Code Llama, or Magicoder), we observe an imbalance in their ability to handle positive and negative cases. While these models can achieve competitive Precision, they struggle with Recall. For example, DeepSeek Coder-6.7B achieves Precision scores of 0.69 and 0.70 in the independent-unit and repo-level code generation settings, respectively, which is comparable to the other approaches. However, its Recall is up to about two times lower than the others. The reason is that these models tends to be over-confident in the correctness of its generated code, frequently labeling its generated code as “correct” even when it is actually incorrect.
In practice, LLM-AJ requires no additional training but exhibits variable performance across settings, as shown by the drop in LLM-AJ’s accuracy between the independent-unit and repo-level code generation scenarios. Meanwhile, Openia delivers consistently robust performance by leveraging a lightweight classifier trained on internal states. To ensure broad generalization, we evaluated Openia across cross-benchmark, cross-task, and repository-level scenarios, each featuring distinct data distributions. In every case, Openia maintained high accuracy, precision, recall, and F1-Score, demonstrating its resilience to domain shifts and varying code structures. For a practical application, practitioners can pre-train Openia on curated data from prior tasks or publicly available datasets. During deployment, new data can be incrementally collected and used to periodically update Openia, thereby improving its performance for task-specific and evolving use cases.
The traditional post-hoct approaches, which predict code correctness using the code semantics encoded by pre-trained models such as CodeBERT and CodeT5/CodeT5+, demonstrate a quite strong performance. Despite their strengths, these baselines are outperformed by Openia, as evidenced by consistently better performance across all models in Table 2 and Table 3. For repo-level code generation, Openia achieves an F1-Score of 0.73 with DeepSeek Coder-6.7B, compared to 0.69 for CodeT5+. Especially for Code Llama-7B, the margin of improvement is even more significant, with Openia attaining F1-Scores of 0.79, respectively, compared to 0.71 and 0.69 for the classifier based on CodeT5+ and CodeBERT. These results suggest that relying on pre-trained models to encode generated code, while effective to an extent, is insufficient for capturing the nuanced, dynamic, and process-oriented aspects of code correctness.
Unlike these black-box (closed-box) approaches that rely solely on final outputs, both Openia and LLM-Check are white-box (open-box) approaches that examines the LLMs’ internal workings to make predictions. LLM-Check combines an Eigen-analysis of hidden states and attention maps with token-level uncertainty quantification to flag hallucinations without any additional training [40]. While this zero-training design makes LLM-Check immediately deployable, its performance on code-quality assessment remains lower than that of Openia up to 40%. By contrast, Openia analyzes the intermediate states of the LLM during code generation and trains a lightweight classifier on these in-process signals. This approach uncovers hidden patterns associated with the generated code, mitigating the overconfidence often seen in LLM-AJ approach and providing a more nuanced and reliable evaluation. Moreover, the internal representations utilized by Openia not only encode the code features as CodeBERT and CodeT5+ but also capture the model’s reasoning process. This reasoning process information reflects not just dataset-specific patterns but also fundamental programming underlying programming principles. Openia can dynamically adapt to varying requirements, programming constructs, and output expectations across different tasks and benchmarks, enhancing its generality and robustness. This adaptability enhances its generality and robustness, enabling it to consistently outperform both post-hoc classifiers and LLM-Check in our experiments across all experimental settings. Additionally, by leveraging these in-process signals, Openia can provide a unique advantage: it can detect potential issues and assess correctness as the code is being generated, enabling real-time quality assurance.
Furthermore, Openia also offers a promising alternative approach for benchmarking LLMs in scenarios where ground-truth test cases are unavailable or costly to produce. Indeed, in our experiments, we found that code samples labeled “correct” by Openia passed actual test cases at rates 50–250% higher than those labeled “incorrect”. Moreover, for each setting, we measured the performance of every studied code LLM in pass@1, as well as the proportions of generated code samples labeled “correct” ($\mathcal{P}$) by Openia. We observed strong positive correlations with the Pearson correlation coefficients between pass@1 and $\mathcal{P}$ are 0.99, 0.73, and 0.97 for Cross-benchmark, Cross-task, and Repo-level settings respectively. These findings demonstrate that models ranked higher by Openia tend to achieve higher test-pass rates in practice. As a result, developers can generate code for identical tasks from different LLMs and use Openia to estimate each model’s relative correctness. The LLM that produces a higher proportion of predicted-correct outputs can be deemed superior. This opens up a practical and test-free solution for LLM comparison in settings where constructing tests is costly.
@@ 5.1.2 Code Length Analysis
Figure 4 illustrates that Openia achieves its best performance on shorter generated code. Specifically, for the code snippets ranging from 1 to 30 tokens, Openia accurately predicts code correctness with an F1-Score of 0.81. This strong performance is attributed to the simplicity of short code snippets, which typically contain few dependencies. Thus, the semantic and syntactic features of the code can be effectively encoded within the LLM’s internal states, thereby facilitating accurate predictions. However, Openia’s F1-score decreases as the code length increases, reaching its lowest point (0.66) for medium-length code (51-70 tokens). The decline in performance is likely due to the increased complexity and dependencies in longer code, which makes it more challenging for the LLM’s internal representations to encode all critical information.
Interestingly, Openia exhibits performance recovery for longer code exceeding 70 tokens, despite the challenges posed by increased complexity. This improvement suggests that Openia can effectively utilize the richer contextual information and redundancy available in the internal states of longer code to reinforce key patterns and enhance correctness predictions. These results highlight the adaptability of Openia in leveraging the LLM’s representations across varying lengths.